OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)) -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (anonymous):

@ganeshie8 @hartnn @mathslover @sleepyjess

OpenStudy (sleepyjess):

Just a minute :)

OpenStudy (sleepyjess):

Sorry, I couldn't load the page.

OpenStudy (sleepyjess):

Ok, all f(g(x)) really means is that you plug in g(x) for x in f(x). I will show you how to do that, then see if you can figure out the next step :) f(x) = -4x + 7 g(x) = 10x - 6 f(g(x)) = -4(10x-6) + 7